# Proportionality for Constrained Public Decisions

### v1

@@ 7 Conclusion
We considered two different interpretations of justified representation from multiwinner voting and adapted them to a public-decision model with constraints. In analysing the feasibility of the axioms, we devised restricted classes of constraints (the NFD property and simple implications). While we could show mostly negative results for the satisfaction of cohesiveness-EJR under constraints, we were able to adapt successfully three known rules (MES, Local Search PAV and MeCorA)
to yield positive proportional guarantees that meet, in an approximate sense, the requirements of agreement-EJR. Additionally, we defined a suitable notion of priceability and showed that our adaptation of MeCorA always returns priceable outcomes.
Our work opens up a variety of paths for future research. First, assessing a class of constraints that are more expressive than the simple implications seems a natural starting point in extending our work. Then, on a more technical level, it would be interesting to check if the representation guarantees that are offered by $\text{MES}_{\mathcal{C}}$, LS-$\text{PAV}_{\mathcal{C}}$ and Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}(k-1)$ still hold for a wider range of election instances. Regarding our adaptation of priceability, the question is open as to whether there are more constrained public-decision that always produce complete priceable outcomes. Given that we opted to represent the constraints as an enumeration of all feasible outcomes, it is natural to ask what occurs to results such as Proposition 1 when we consider the constraint takes a particular form of representation, e.g., $\mathcal{C}$ is represented as a Boolean formula of propositional logic. We also note some lingering computational questions such as the computational complexity of $(i)$ computing outcomes for rules such as $\text{MES}_{\mathcal{C}}$ and Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}(k-1)$ for general constraints, and $(ii)$ of checking whether a given feasible outcome is priceable. Finally, the list of proportionality notions to be tested on the constraints test-bed is not exhausted, with the proportionality degree (Lackner and Skowron, 2023) most notably still to be considered.

### v2

@@ 7 Conclusion
We considered two different interpretations of justified representation from committee elections and adapted them to a public-decision model with constraints. In analysing the feasibility of the axioms, we devised restricted classes of constraints such as those not fixing decisions and those that can be represented as simple implications. While we could show mostly negative results for the satisfaction of cohesiveness-EJR under constraints, we were able to adapt successfully three known rules (MES, Local Search PAV and MeCorA) to yield positive proportional guarantees that meet, in an approximate sense, the requirements of agreement-EJR. Additionally, we defined a suitable notion of priceability and showed that our adaptation of MeCorA always returns priceable outcomes.
Our work opens up a variety of paths for future research. First, assessing a class of constraints that are more expressive than simple implications seems a natural starting point in extending our work. Then, on a more technical level, it would be interesting to check if the representation guarantees that are offered by $\text{MES}_{\mathcal{C}}$, LS-$\text{PAV}_{\mathcal{C}}$ and Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}(k-1)$ still hold for a wider range of election instances. Regarding our adaptation of priceability, the question is open as to whether there are more constrained public-decision rules that always produce complete priceable outcomes. Given that we opted to represent the constraints as an enumeration of all feasible outcomes, it is natural to ask whether our computational results still hold under compact constraint representations, e.g., $\mathcal{C}$ is represented as a Boolean formula of propositional logic. We also note some lingering computational questions such as the computational complexity of $(i)$ computing outcomes for rules such as $\text{MES}_{\mathcal{C}}$ and Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}(k-1)$ for general constraints, and $(ii)$ of checking whether a given feasible outcome is priceable. Finally, the list of proportionality notions to be tested on the constraints test-bed is not exhausted (Lackner and Skowron, 2023), with the proportionality degree (Skowron, 2021) and $\text{EJR}{+}$ (Brill and Peters, 2023) being notable notions still to be considered.